A plane can fly 2250 miles in the same time as it takes a car to go 360 miles. If the car travels 315 mph slower than the plane, find the speed of the plane.
Answer by richwmiller(17219)

m/mph=2250/x=360/(x-315)
x = 375 speed of plane
2250/375=360/60
6 hours traveling time